November 1997  ICTCM X (Chicago)  3

Computer Algebra Tools for Solving BVPs

> ODE := diff( y(x), x$2 ) + y(x) = 0:
> BCa := y(0)=2, y(Pi/2)=0:
> SOLa := dsolve( { ODE, BCa }, y(x) );

SOLa := y(x) = 2 cos(x)

> BCb := y(0)=2, y(Pi)=0:
> SOLb := dsolve( { ODE, BCb }, y(x) );

SOLb :=

> BCc := y(0)=2, y(Pi)=-2:
> SOLc := dsolve( { ODE, BCc }, y(x) );

SOLc := y(x) = C1 sin(x) + 2 cos(x)